OpenStudy (anonymous):

Solve for x. -3 + 8x - 5 = -8

OpenStudy (anonymous):

collect like terms

OpenStudy (anonymous):

it would be more beneficial to learn the steps, instead of people just writing the answers

OpenStudy (nurali):

-3 + 8x - 5 = -8 8x=-8+5+3 8x=-8+8 8x=0 x=0/8 x=0